OpenStudy (anonymous):

3x=√30-3x A. −2 B. −5/3 C. 5/3 and −2 D. 5/3

OpenStudy (studygurl14):

Is it \[3x = \sqrt{30} -3x\] or \[3x = \sqrt{30 - 3x}\] ?

OpenStudy (anonymous):

The 2nd one @StudyGurl14

OpenStudy (studygurl14):

Okay. So, to solve this, all you have to do is square both sides ot get \[9x ^{2} = 30 - 3x\] Can you do the rest?

OpenStudy (anonymous):

Yes, I can do the rest. My teacher confused me with this part because he messed up on the step when he was teaching this.. thank you so much!!

OpenStudy (studygurl14):

You're welcome. your teacher messed up? lol

OpenStudy (anonymous):

Yes he did lol. He does that quite often!